OpenStudy (anonymous):

Lee earns $6.25 per hour mowing lawns. a. Write a rule to describe how the amount of money m earned is a function of the number of hours h spent mowing lawns. b. How much does Lee earn if she works 5 hours and 15 minutes?

OpenStudy (the_fizicx99):

f(m) = 6.25m where is the how many hours she works

OpenStudy (the_fizicx99):

6.25*5 = 31.25 1/4 of an hour is 15 min 1/4 = .25 * 4 = 1.56 31.25 + 1.56 = 32.81